# Planets and planetesimals at cosmic dawn: Vortices as planetary nurseries

### v1

@@ 1 Introduction
Because collisional growth beyond pebble size is hindered by the previously mentioned barriers, an alternative mechanism must facilitate the formation of larger bodies. This mechanism is gravitational collapse, which leads to the formation of $\sim 1-1000\,\textrm{km}$-sized planetesimals. For gravitational collapse to occur, the solids density must reach the so-called Roche density, which is significantly higher than the typical solids density in the disk. One favored mechanism for concentrating dust is the streaming instability (SI, Youdin & Goodman 2005; Johansen & Youdin 2007), which arises due to the mutual drag between solids and gas in the disk and leads to the formation of dense filaments that can locally reach Roche density. Much work has gone into identifying the conditions under which the SI leads to planetesimal formation (Carrera et al., 2015; Yang et al., 2017; Li & Youdin, 2021; Lim et al., 2024a). Based on these criteria, the SI would not operate in low-metallicity disks such as the ones considered in this work unless a separate mechanism first acts to concentrate the dust.
@@ 2 Numerical method
We consider a viscously evolving protoplanetary disk and simulate the growth and transport of dust within it. The effect of vortices is modeled via post-processing, and we compare our results with SI criteria to determine when and where planetesimals form. We then simulate the further growth of these planetesimals via pebble accretion, accounting for the enhanced accretion rate inside vortices.
@@ 2.1 Evolution of the gas and dust disk
where $k_{\rm B}$ is the Bolzmann constant, $T$ is the mid-plane temperature, $\mu$ is the mean molecular weight and $m_{\rm H}$ is the mass of the hydrogen atom. The temperature profile is taken to be that of a passively irradiated disk with an irradiation angle of $0.05$:
@@
We motivate our initial conditions from the primordial star formation simulations of Sharda & Menon (2024) – specifically of a low-mass star that formed through fragmentation within the disc of a massive protostar, and is ejected from the disk via many-body interactions, subsequently evolving quasi-independently as an isolated star-disk system. We adopt its stellar mass of $0.45\,\textrm{M}_{\odot}$ and its corresponding initial disk mass of $0.0071\,\textrm{M}_{\odot}$ in our main study, which corresponds to an initial disk accretion rate of $\dot{M}_{0}=1.32\times 10^{-8}\,\textrm{M}_{\odot}\textrm{yr}^{-1}$ (calculated using a disk ranging from $1-200\,\textrm{au}$). We adopt stellar radii and effective temperatures from Baraffe et al. (2015) ($R_{*}=2.411\,\textrm{R}_{\odot}$ and $T_{*}=3761\,\textrm{K}$ for a $0.45\,\textrm{M}_{\odot}$ star). In the Appendix, we also present results from simulations performed with a solar mass star and a disk-to-star mass ratio of $0.1$ and $0.01$.
@@ 2.2 Concentration and growth of dust in vortices
In this section, we describe how the DustPy data is post-processed to obtain what the dust density and $\mathrm{St}$ would have been if a vortex was present in the disk. This post-processing is applied to every point on the time and semimajor axis grid. We consider two distinct scenarios for dust concentration and growth inside the vortex. Case VC for "vortex concentration" represents a scenario where dust is concentrated inside the vortex, but there is no additional dust growth compared to the unperturbed disk. Case VCG for "vortex concentration plus growth" represents a scenario where both concentration and growth are promoted inside the vortex, leading to a feedback loop due to the interdependence of these processes.
@@
In this expression, $\rho_{\rm dust}$ and $\rho_{\rm gas}$ are the dust and gas density, respectively, and $Z$ is the dust-to-gas surface density ratio of the unperturbed disk. The above expression is derived for a mono-disperse dust size distribution. Since DustPy outputs an entire size distribution, we need to choose a representative $\mathrm{St}$, which we pick to be the density-weighted average $\mathrm{St}$. In Case VC, we post-process the DustPy data using Eq. 6 and this representative $\mathrm{St}$ to obtain the final dust-to-gas ratios. In this model, we do not consider the time required to reach $\epsilon_{\rm max}$, but assume that it occurs instantaneously.
In Case VCG, we account for both the concentration and growth of dust inside the vortex, while also considering the effects of mass loading. The full derivation and implementation of this case are provided in Carrera et al. (in prep); here, we present a summary. Studies of dust growth inside vortices and pressure bumps find that grains grow to significantly larger sizes compared to the background disk (Bae et al., 2018; Li et al., 2020). Due to the lack of fast radial drift, it is reasonable to assume that growth proceeds until the fragmentation limit. When accounting for mass loading, we find that the expression for $\mathrm{St}_{\rm frag}$ depends positively on the dust density. Similarly, the expression for the peak dust density has a positive dependence on $\mathrm{St}$ (see Eq. 6). As a result, when dust concentrates at the center of the vortex, grains can grow to larger sizes, leading to even higher concentrations, creating a positive feedback loop. We note that, while theoretically motivated, the described feedback loop has yet to be numerically tested.
@@
Which of these end states is reached depends on the fragmentation barrier, strength of turbulence, as well as the relative timescales of dust growth $t_{\mathrm{St}}$ and vortex concentration $t_{Z}$. Birnstiel et al. (2012) derived the dust growth timescale in a protoplanetary disk as $t_{\mathrm{St}}=1/(Z\Omega)$. Here we follow Carrera et al. (in prep) and assume that, since a vortex is in many ways analogous to a small protoplanetary disk, the dust growth timescale in a vortex is $t_{\mathrm{St}}=1/(Z\Omega_{\mathrm{vortex}})$, where $\Omega_{\mathrm{vortex}}$ is the vortex angular frequency. In practice, this has no meaningful effect on the simulation results since the two frequencies are almost identical. Following Lyra & Lin (2013) and Carrera et al. (in prep) we adopt $\Omega_{\mathrm{vortex}}=0.5\Omega$. Finally, Carrera et al. (in prep) derived the vortex concentration timescale $t_{Z}=1/(\mathrm{St}\Omega_{\mathrm{vortex}})$.
We assume that the vortex has a finite lifetime, which sets an upper limit on the time available for growth and concentration to proceed. We adopt a vortex lifetime of 1000 orbital periods in our main study. This is motivated by the re-generation of the m=1 mode of vortices from Rossby-Wave-Instability cycles found in Regály et al. (2012), which was found to be of the order $\sim$1000 orbits. We discuss how shorter or longer vortex lifetimes affect our results in Section 5.2.2. To keep the model simple and conservative, we assume that the vortex’s total dust budget remains constant after formation; in other words, we do not account for pebbles drifting into the vortex post-formation. Additional information regarding the derivation and implementation of case VCG are shown in Sect. D.
@@ 2.3 Planetesimal formation
The conditions under which the SI leads to planetesimal formation remains an active area of research. Perhaps the most reliable SI criterion comes from Lim et al. (2024b), who conducted a large number of 3D simulations of the SI under external (i.e., forced) turbulence. This represents a significant improvement over earlier criteria, which were based on 2D simulations without external turbulence. We summarize their criterion here, which we refer to as SIa:
@@
Importantly, Lim et al. (2024b) showed that expressing the criterion in terms of $\epsilon$ instead of $Z$ makes the criterion remain valid for a range of alpha values (they tested $10^{-4}\leq\alpha_{\rm t}\leq 10^{-3}$). The one limitation of this work is that, due to the cost of 3D simulations, they could only model a limited range of Stokes numbers $(0.01\leq\mathrm{St}\leq 0.1)$, while grains in in the disk can be much smaller than this range.
To address this limitation, we also consider the criterion by Lim et al. (2024a) (henceforth referred to as SIb), which is currently the most comprehensive among the 2D criteria available in the literature:
@@
(Liu et al., 2020; Lorek & Johansen, 2022). In the above equation $\zeta=4\pi G\rho_{\rm gas}\Omega_{\rm K}^{-2}$ is the self-gravity parameter, $h=H/r$ is the disk aspect ratio and $Z_{\rm fil}$ is the filament dust-to-gas ratio, here taken to be $0.1$. The planetesimals most likely to continue growing into planets are those at the upper end of the size distribution. We follow Lyra et al. (2023), who assumed that the most massive planetesimal is $\sim$10x more massive than the typical planetesimal mass. We refer to this object as an embryo, and further limit the embryo mass by the total mass of dust trapped within the vortex.
@@ 2.4 Pebble accretion
How much the embryo grows by pebble accretion is strongly influenced by the residence time of the embryo within the vortex (shown in Fig. 6). Since we are considering very low dust-to-gas ratios in this work, the criteria for forming planetesimals can only be reached towards the vortex center, where the dust density is high and the residence time is long. We choose to adopt a nominal residence time of 10 orbital periods.
@@ 3 Formation of the first planetesimals
Figure 1: The plot shows the Stokes number (top row), midplane dust-to-gas density ratio (middle row) and planetesimal formation regions along with the resulting embryo masses (bottom row) for our simulation with $v_{\rm frag}=1\,\textrm{m}\,\textrm{s}^{-1}$, $Z=1.6\times 10^{-3}$ and $\alpha_{\rm T}=5\times 10^{-5}$. When no vortices are assumed (left column), the dust-to-gas ratios are much too low to trigger planetesimal formation. When vortices are present and act to concentrate the dust without affecting dust growth (middle column), the peak dust-to-gas ratios increase significantly, leading to large regions of planetesimal formation. When both concentration and growth are enhanced inside the vortex (right column), planetesimals can form almost everywhere.
Figure 2: The plots show the regions of planetesimal formation along with the corresponding embryo masses for all simulations in the parameter study. The colored lines outline the regions where planetesimals form in case VC and VCG, for both criteria SIa and SIb.
@@
The middle column of Fig. 1 shows the maximum $\epsilon$ that would have been obtained if vortices were present that concentrated the dust (case VC). The bottom plot shows where in time and space the planetesimal formation criteria are met (here we use Eq.10 for the SI criteria), along with the masses of the formed embryos. It is important to note that, although our post-processing is applied to every point on the space-time grid, the correct interpretation of the plot is as follows: if a vortex were to form at a specific time and location, the corresponding values of $\mathrm{St}$ and $\epsilon$ would be obtained at the vortex center, indicating whether planetesimal formation could occur. In case VC, the maximum dust-to-gas midplane density ratio obtained at the vortex center exceeds the SI criteria in large parts of the time-space grid. Due to the low mass of the star and disk, the masses of the forming embryos are relatively small, on the order $10^{-5}\,\textrm{M}_{\oplus}$.
@@
The right column of Fig. 1 presents the results for case VCG, where both concentration and growth are enhanced inside the vortex, resulting in a growth-concentration feedback loop due to the interdependence of these processes. In this scenario, the criteria for forming planetesimals were reached almost everywhere on the space-time grid. The black contour in the $\mathrm{St}$ plot outlines the region where $\mathrm{St}\geq 1$ is reached before the Roche density.
@@
Fig. 2 shows the regions of planetesimal formation along with the corresponding embryo masses for all simulations in the parameter study. In the scenario where dust sizes remain the same as in the unperturbed disk and concentration occurs instantaneously (case VC), vortices can trigger planetesimal formation for metallicities as low as $Z=0.0004$ if $\alpha_{\rm t}=10^{-5}$. For $\alpha_{\rm t}=5\times 10^{-5}$, a dust-to-gas ratio of $Z=0.0008$ or $Z=0.0016$ is required, depending on the assumed fragmentation velocity and SI criteria. At higher turbulence levels, merely concentrating dust inside vortices is no longer sufficient to trigger planetesimal formation. The exact region of planetesimal formation depends on the assumed SI criteria. Criterion SIb allows for planetesimal formation in a wider region than criterion SIa, which is expected since criterion SIa accounts for the effects of turbulence.
@@
In the scenario where dust growth and concentration boost each other in a feedback cycle (case VCG), planetesimal formation inside vortices is also possible for $\alpha_{\rm t}>5\times 10^{-5}$. When considering a fragmentation velocity of $1\,\textrm{ms}^{-1}$, planetesimal formation in vortices is possible for $\alpha_{\rm t}\leq 10^{-4}$. Using a fragmentation velocity ten times larger allows for planetesimal formation even when $\alpha_{\rm t}>10^{-4}$, provided $Z\gtrsim 0.0008$. Increasing the fragmentation velocity beyond $1\,\textrm{ms}^{-1}$ has little effect when $\alpha_{\rm t}<10^{-4}$, since particle growth in this case is mostly limited by radial drift.
@@
When considering $\alpha_{\rm t}=10^{-5}$ and SI criterion SIb, case VCG results in narrower planetesimal formation regions than case VC. This is because case VCG accounts for the growth and concentration timescales, whereas case VC assumes instantaneous concentration. Furthermore, the results of case VCG are highly dependent on the assumed vortex lifetime. A longer vortex lifetime - or equivalently, shorter growth and concentration timescales - would allow for additional dust growth and concentration. Fig. 7 shows how the planetesimal formation regions in case VCG change when assuming a vortex lifetime that is 10 times longer. In this scenario, planetesimal formation becomes possible even for $Z=0.0002$.

### v2

@@ 1 Introduction
[…]
@@ 2 Numerical method
We consider a viscously evolving protoplanetary disk and simulate the growth and transport of dust within it. The effect of vortices is modeled via post-processing, and we compare our results with SI criteria to determine when and where planetesimals form. We then simulate the further growth of these planetesimals via pebble accretion, accounting for the enhanced accretion rate inside vortices. The code used and developed for this work is open-source and available on github [^1] https://github.com/astrolinn/firstPlanets.
@@ 2.1 Evolution of the gas and dust disk
where $k_{\rm B}$ is the Bolzmann constant, $T$ is the mid-plane temperature, $\mu$ is the mean molecular weight and $m_{\rm H}$ is the mass of the hydrogen atom. The temperature profile is taken to be that of a passively irradiated disk with an irradiation angle of 0.050.050.05:
@@
[…]
@@ 2.2 Concentration and growth of dust in vortices
In this section, we describe how the DustPy data is post-processed to obtain what the dust density and $\mathrm{St}$ would have been in the presence of vortices. We consider two distinct scenarios for dust concentration and growth inside a vortex. Case VC for "vortex concentration" represents a scenario where dust is concentrated inside the vortex, but there is no additional dust growth compared to the unperturbed disk. Case VCG for "vortex concentration plus growth" represents a scenario where both concentration and growth are promoted inside the vortex, leading to a feedback loop due to the interdependence of these processes.
@@
This is a drift-diffusion equilibrium model, in which $\rho_{\rm dust}$ and $\rho_{\rm gas}$ are the dust and gas midplane density, respectively, and $Z$ is the dust-to-gas surface density ratio of the unperturbed disk. The above expression is derived for a mono-disperse dust size distribution. Since DustPy outputs an entire size distribution, we need to choose a representative $\mathrm{St}$, which we pick to be the density-weighted average $\mathrm{St}$.
In Case VC, we post-process the DustPy data using Eq. 6 to obtain what the peak dust-to-gas midplane density ratio would be at each point on the time and semimajor axis grid if a vortex were to form there. In this model, we do not consider the time required to form the vortex or the time needed to reach drift-diffusion equilibrium, but assume that it occurs instantaneously. Since we assume that the $\mathrm{St}$ remain the same as in the unperturbed disk - i.e., no growth or destruction of dust inside the vortex - there is no dust coagulation timescale involved.
In Case VCG, we relax the last assumption and account for both the concentration and growth of dust inside the vortex, while also considering the effects of mass loading. The full derivation and implementation of this case are provided in Carrera et al. (2025); here, we present a summary. Studies of dust growth inside vortices and pressure bumps find that grains grow to significantly larger sizes compared to the background disk (Bae et al., 2018; Li et al., 2020). Due to the lack of fast radial drift, it is reasonable to assume that growth proceeds until the fragmentation limit. When accounting for mass loading, we find that the expression for $\mathrm{St}_{\rm frag}$ depends positively on the dust density. Similarly, the expression for the peak dust density has a positive dependence on $\mathrm{St}$ (see Eq. 6). As a result, when dust concentrates at the center of the vortex, grains can grow to larger sizes, leading to even higher concentrations, creating a positive feedback loop. We note that, while theoretically motivated, the described feedback loop has yet to be numerically tested.
@@
Which of these end states is reached depends on the fragmentation barrier, strength of turbulence, as well as the relative timescales of dust growth $t_{\mathrm{St}}$ and vortex concentration $t_{Z}$. Birnstiel et al. (2012) derived the dust growth timescale in a protoplanetary disk as $t_{\mathrm{St}}=1/(Z\Omega)$. Here we follow Carrera et al. (2025) and assume that, since a vortex is in many ways analogous to a small protoplanetary disk, the dust growth timescale in a vortex is $t_{\mathrm{St}}=1/(Z\Omega_{\mathrm{vortex}})$, where $\Omega_{\mathrm{vortex}}$ is the vortex angular frequency. In practice, this has no meaningful effect on the simulation results since the two frequencies are almost identical. Following Lyra & Lin (2013) and Carrera et al. (2025) we adopt $\Omega_{\mathrm{vortex}}=0.5\Omega$. Finally, Carrera et al. (2025) derived the vortex concentration timescale $t_{Z}=1/(\mathrm{St}\Omega_{\mathrm{vortex}})$. Grain growth and vortex trapping occur in tandem, each on its respective timescale. For example, if drift-diffusion occurs on a shorter timescale, dust grains still grow toward $\mathrm{St}_{\rm frag}$, but not as quickly as the concentration moves toward drift-diffusion equilibrium. In this scenario, $\mathrm{St}_{\rm frag}$ is essentially a moving target for dust growth. We follow Carrera et al. (2025) and set the iteration timestep to $dt=\min(t_{\mathrm{St}},t_{Z})$, allowing the two processes to evolve gradually within a changing environment.
We assume that the vortex has a finite lifetime, which sets an upper limit on the time available for growth and concentration to proceed. We adopt a vortex lifetime of 1000 orbital periods in our main study. This is motivated by the re-generation of the m=1 mode of vortices from Rossby-Wave-Instability cycles found in Regály et al. (2012), which was found to be of the order $\sim$1000 orbits. We discuss how shorter or longer vortex lifetimes affect our results in Section 5.2.2. To keep the model simple and conservative, we assume that the vortex’s total dust budget remains constant after formation; in other words, we do not account for pebbles drifting into the vortex post-formation.
We perform the post-processing at every point on the time and semimajor axis grid. The results should be interpreted as follows: for a vortex formed at $(r=r_{j},t=t_{i})$, the model outcomes $\mathrm{St}(j,i)$ and $\epsilon_{\rm max}(j,i)$ are the maximum Stokes number and dust-to-gas midplane density ratio that would be obtained at the vortex center during its 1000-orbit lifetime. Additional information regarding the derivation and implementation of case VCG is shown in Sect. D.
@@ 2.3 Planetesimal formation
The conditions under which the SI leads to planetesimal formation remains an active area of research. Perhaps the most reliable SI criterion comes from Lim et al. (2024), who conducted a large number of 3D simulations of the SI under external (i.e., forced) turbulence. This represents a significant improvement over earlier criteria, which were based on 2D simulations without external turbulence. We summarize their criterion here, which we refer to as SI24:
@@
Importantly, Lim et al. (2024) showed that expressing the criterion in terms of $\epsilon$ instead of $Z$ makes the criterion remain valid for a range of alpha values (they tested $10^{-4}\leq\alpha_{\rm t}\leq 10^{-3}$). The one limitation of this work is that, due to the cost of 3D simulations, they could only model a limited range of Stokes numbers $(0.01\leq\mathrm{St}\leq 0.1)$, while grains in in the disk can be much smaller than this range.
To address this limitation, we also consider the criterion by Lim et al. (2025) (henceforth referred to as SI25), which is currently the most comprehensive among the 2D criteria available in the literature:
@@
(Liu et al., 2020; Lorek & Johansen, 2022). In the above equation $\zeta=4\pi G\rho_{\rm gas}\Omega_{\rm K}^{-2}$ is the self-gravity parameter, $h=H/r$ is the disk aspect ratio and $Z_{\rm fil}$ is the filament dust-to-gas ratio, here taken to be 0.10.10.1. The planetesimals most likely to continue growing into planets are those at the upper end of the size distribution. We follow Lyra et al. (2023), who assumed that the most massive planetesimal is $\sim$10x more massive than the typical planetesimal mass. We refer to this object as an embryo, and further limit the embryo mass by the total mass of dust trapped within the vortex.
@@ 2.4 Pebble accretion
How much the embryo grows by pebble accretion is strongly influenced by the residence time of the embryo within the vortex (shown in Fig. 6). Since we are considering very low dust-to-gas ratios in this work, the criteria for forming planetesimals can only be reached near the vortex center, where the dust density is high and the residence time is long. We adopt a nominal residence time of 10 orbital periods. We make the simplified assumption that each vortex forms only one embryo, and the remainder of the solid budget remains in dust and pebbles (see Sect. 5.2.4 for a discussion of this assumption).
@@ 3 Formation of the first planetesimals
Figure 1: The plot shows the representative Stokes number (top row), midplane dust-to-gas density ratio $\epsilon$ ($\epsilon_{\rm max}$ for case VC/VCG, middle row), and planetesimal formation regions along with the resulting embryo masses (bottom row) for our simulation with $v_{\rm frag}=1\,\textrm{m}\,\textrm{s}^{-1}$, $Z=1.6\times 10^{-3}$, and $\alpha_{\rm T}=5\times 10^{-5}$. If there are no vortices (left column), $\epsilon$ is much too low to trigger planetesimal formation. The middle column shows the $\epsilon_{\rm max}$ and $\mathrm{St}$ that would be obtained at the center of a vortex formed at each individual space-time coordinate ($r_{j},t_{i}$), assuming vortices act to concentrate the dust without affecting dust growth. In this scenario, $\epsilon_{\rm max}$ is significantly enhanced compared to the unperturbed disk, and vortex formation would result in planetesimal formation in large parts of the parameter space. If both concentration and growth are enhanced inside the vortex (right column), vortex formation would lead to planetesimal formation almost regardless of when are where the vortex forms.
Figure 2: The plots show where on the space-time grid vortex formation would result in planetesimal formation for all simulations in the parameter study, along with the corresponding embryo masses. The colored lines outline the regions where planetesimals form in case VC and VCG, for both criteria SI24 and SI25.
@@
The middle column of Fig. 1 shows the $\mathrm{St}$ and peak dust-to-gas midplane density ratio $\epsilon_{\rm max}$ that would be obtained at the center of a vortex formed at each individual space-time coordinate ($r_{j},t_{i}$), if vortices act to concentrate the dust without affecting dust growth (case VC). The bottom plot shows where in time and semimajor axis space vortex formation would result in planetesimal formation (here we use Eq.10 for the SI criteria), along with the masses of the formed embryos. In case VC, $\epsilon_{\rm max}$ exceeds the SI criterion in large parts of the parameter space. Due to the low mass of the star and disk, the masses of the forming embryos are relatively small, on the order of $10^{-5}\,\textrm{M}_{\oplus}$.
@@
The right column of Fig. 1 presents the results for case VCG, where both concentration and growth are enhanced inside the vortex, resulting in a growth-concentration feedback loop due to the interdependence of these processes. The presented $\mathrm{St}$ and $\epsilon_{\rm max}$ are the maximum Stokes number and dust-to-gas midplane density ratio that would be obtained at the center of a vortex during its 1000-orbit lifetime. In this scenario, vortex formation would result in planetesimal formation almost everywhere on the space-time grid, except at large semimajor axes when the disk is still young. This is because the dust densities and sizes are too small at the time of vortex formation, resulting in long growth and concentration timescales inside the vortex. The sharp transition between the region where the feedback loop is efficient and inefficient is highly dependent on the assumed vortex lifetime (see Fig. 7). The black contour in the $\mathrm{St}$ plot outlines the region where $\mathrm{St}\geq 1$ is reached before the Roche density.
@@
Fig. 2 shows where on the space-time grid vortex formation would result in planetesimal formation for all simulations in the parameter study, along with the corresponding embryo masses. In the scenario where dust sizes remain the same as in the unperturbed disk and concentration occurs instantaneously (case VC), vortices can trigger planetesimal formation for metallicities as low as $Z=0.0004$ if $\alpha_{\rm t}=10^{-5}$. For $\alpha_{\rm t}=5\times 10^{-5}$, a dust-to-gas ratio of $Z=0.0008$ or $Z=0.0016$ is required, depending on the assumed fragmentation velocity and SI criteria. At higher turbulence levels, merely concentrating dust inside vortices is no longer sufficient to trigger planetesimal formation. The exact space-time region where vortex formation would result in planetesimal formation depends on the assumed SI criteria. Criterion SI25 allows for planetesimal formation inside vortices in a wider region of parameter space than criterion SI24, which is expected, since criterion SI24 accounts for the effects of turbulence.
@@
In the scenario where dust growth and concentration boost each other in a feedback cycle (case VCG), planetesimal formation inside vortices is also possible for $\alpha_{\rm t}>5\times 10^{-5}$. When considering a fragmentation velocity of $1\,\textrm{ms}^{-1}$, planetesimal formation inside vortices is possible for $\alpha_{\rm t}\leq 10^{-4}$. Using a fragmentation velocity ten times larger allows for planetesimal formation even when $\alpha_{\rm t}>10^{-4}$, provided $Z\gtrsim 0.0008$. Increasing the fragmentation velocity beyond $1\,\textrm{ms}^{-1}$ has little effect when $\alpha_{\rm t}<10^{-4}$, since particle growth in this case is mostly limited by radial drift.
@@
When considering $\alpha_{\rm t}=10^{-5}$ and SI criterion SI25, case VCG results in narrower planetesimal formation regions than case VC. This is because case VCG accounts for the growth and concentration timescales inside the vortex, whereas case VC assumes that drift-diffusion equilibrium is reached instantaneously. Furthermore, the results of case VCG are highly dependent on the assumed vortex lifetime. A longer vortex lifetime - or equivalently, shorter growth and concentration timescales - would allow for additional dust growth and concentration. Fig. 7 shows how the region where vortex formation results in planetesimal formation changes when assuming a vortex lifetime that is 10 times longer. In this scenario, planetesimal formation becomes possible even for $Z=0.0002$.